# Orient Anything

### v1

@@ 1 Introduction
Orientation estimation is a fundamental task in 3D shape analysis which consists of estimating a shape’s orientation axes: its side-, up-, and front-axes. Using this data, one can rotate a shape into canonical orientation, in which the shape’s orientation axes are aligned with the coordinate axes. This task is especially important as a pre-processing step in 3D deep learning, where deep networks are typically trained on datasets of canonically-oriented shapes but applied to arbitrarily-oriented shapes at inference time. While data augmentation or equivariant and invariant architectures may improve a model’s robustness to input rotations, these techniques come at the cost of data efficiency and model expressivity (Kuchnik & Smith, 2019; Kim et al., 2023). In contrast, orientation estimation allows one to pre-process shapes at inference time so that their orientation matches a model’s training data.
Orientation estimation is a challenging task, and developing an orientation pipeline that reliably estimates complete orientations of general shapes remains an open problem. The naïve deep learning approach is to train a model with an $L^{2}$ loss to directly predict a shape’s orientation from a point cloud of surface samples. However, this strategy fails for shapes with rotational symmetries, where the optimal solution to the $L^{2}$ regression problem is the Euclidean mean (Moakher, 2002) of a shape’s orientations over all of its symmetries. In contrast, works such as Poursaeed et al. (2020) discretize the unit sphere into a set of fixed rotations and train a classifier to predict a probability distribution over these rotations, but find that this approach fails for any sufficiently dense discretization of the unit sphere.
@@
Using this strategy, our method achieves state-of-the-art performance on the well-studied problem of up-axis prediction, and additionally performs well on full-orientation prediction, which few prior works have tackled. Unlike previous work, we train and evaluate our model on the entire Shapenet dataset rather than a subset of classes. We further demonstrate its generalization capabilities on Objaverse, a large dataset of real-world 3D models of varying quality.
@@
Our contributions include the following: (1) we identify fundamental obstacles to orientation estimation and study the conditions under which a naïve regression-based approach to orientation estimation fails; (2) we propose a two-stage orientation estimation pipeline that sidesteps these obstacles; (3) we train and test our model on Shapenet and show that it achieves SOTA performance for orientation estimation; (4) we use conformal prediction to enable end users to resolve ambiguities in a shape’s orientation; (5) we release our code and model weights to share our work with the ML community.
@@ 2 Related work
Chen et al. (2021) use reinforcement learning to train a model to gradually rotate a shape into upright orientation. While this algorithm performs well, it is evaluated on few classes and is costly to train. Kim et al. (2020) adopt a similar perspective to Fu et al. (2008), but use ConvNets to extract features for a random forest classifier that predicts a natural base. Poursaeed et al. (2020) use orientation estimation as a pretext task to learn features for shape classification and keypoint prediction.
We also highlight a related literature on canonical alignment. This literature includes works such as Kim et al. (2023); Sajnani et al. (2022); Spezialetti et al. (2020), which seek to map arbitrarily-rotated shapes to a class-consistent pose, as well as Katzir et al. (2022); Sun et al. (2021), which seek to learn pose-invariant representations of 3D shapes. These works only attempt to learn a consistent orientation within each class, but this orientation is not consistent across classes and is not generally aligned with the coordinate axes. In contrast, we tackle the more challenging task of inferring a canonical orientation that is consistent across all objects.
@@ 3 Method
In this section, we motivate and describe our orientation pipeline. We first identify fundamental obstacles to orientation estimation and show that learning a shape’s orientation with the $L^{2}$ loss fails when the shape is rotationally symmetric. Motivated by these observations, we introduce our two-stage orientation pipeline consisting of a quotient orienter followed by a flipper. Our quotient orienter model solves a regression problem to recover a shape’s orientation up to octahedral symmetries, which commonly occur in real-world shapes. The flipper then predicts one of 24 octahedral flips that returns the first-stage output to canonical orientation. We finally use conformal prediction to enable our flipper to output prediction sets whose size varies with the model’s uncertainty. This allows end users to resolve ambiguities in a shape’s orientation by choosing from a small set of plausible candidate orientations.
@@ 3.1 Orientation estimation under rotational symmetries
In this section, we introduce the orientation estimation problem and motivate our approach. Throughout these preliminaries, we consider 3D shapes $S\in\mathcal{S}$ lying in some space of arbitrary shape representations $\mathcal{S}$. Orientation estimation consists of learning an orienter function $f:\mathcal{S}\rightarrow SO(3)$ that maps a shape $S\in\mathcal{S}$ to a predicted orientation $\hat{\Omega}_{S}\in SO(3)$, where $SO(3)$ denotes the 3D rotation group. An orientation is a rotation matrix $\Omega_{S}$ associated with a shape $S$ that is rotation-equivariant: If one rotates $S$ by $R\in SO(3)$ to obtain $RS$, then $\Omega_{RS}=R\Omega_{S}$. We interpret the columns of $\Omega_{S}=(\omega_{S}^{x},\omega_{S}^{y},\omega_{S}^{z})$ as the side-, up-, and front-axes of $S$, respectively, and say that $S$ is in canonical orientation if $\Omega_{S}=I$. If $S$ is in canonical orientation, then its side-, up-, and front-axes are aligned with the $\{x,y,z\}$ coordinate axes, respectively. We depict a canonically-oriented shape $S$ along with its orientation $\Omega_{S}$ in Figure 2.
Given a training set $\mathcal{D}$ of shapes $S\in\mathcal{S}$ paired with their ground truth orientations $\Omega_{S}$, a natural strategy for orientation estimation is to define a loss function $\ell$ on the space of orientations, parametrize $f$ as a neural network $f_{\theta}$, and solve the following problem:
$\min_{f_{\theta}}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\ell\left(f_{\theta}(RS),R,\Omega_{S}\right)\right],$ (1)
where $U(SO(3))$ is the uniform distribution over $SO(3)$. In the following section, we will motivate our approach by describing a theoretical obstacle to learning an orienter function $f$ regardless of one’s choice of loss $\ell$, and then describe specific challenges associated with solving Equation 1 using the $L^{2}$ loss.
Theoretical challenges.
An ideal orienter $f$ should satisfy two desiderata: (1) It should accept arbitrarily-oriented shapes $RS$ as input and output their orientation $\Omega_{RS}=R\Omega_{S}$, and (2) it should succeed on most shapes $S$ that occur in the wild. As many real-world shapes possess at least one non-trivial rotational symmetry, an ideal orienter should therefore succeed on rotationally-symmetric shapes. We begin by showing that no function can simultaneously satisfy these desiderata.
Proposition 3.1
We prove this result in Appendix A.1. Intuitively, if a shape $S$ is symmetric under some non-trivial rotation $R$, then $S$ is invariant under $R$, but its orientation is equivariant under $R$, so the map $RS\mapsto R\Omega_{S}$ is one-to-many and cannot be a function. This shows that any solution to the orientation estimation problem 1 necessarily trades off some desirable property; there are no functions that can successfully orient a rotationally-symmetric shape given any input orientation.
Previous works such as Liu et al. (2016); Poursaeed et al. (2020) observe that orientation estimation via $L^{2}$ regression typically yields poor results. Motivated by this observation, we now characterize the solution to orientation estimation via $L^{2}$ regression for a single rotationally-symmetric shape and show that naïve $L^{2}$ regression degenerates in this setting.
Proposition 3.2
@@
We prove this proposition in Appendix A.2. Proposition 3.2 shows that even when seeking to predict the rotated orientations $R\Omega_{S}$ of a single rotationally-symmetric shape $S$, $L^{2}$ regression fails to learn the correct solution, and instead learns the Euclidean mean of the rotated orientations $RQ\Omega_{S}$ across all rotations $Q$ in the symmetry group $\mathcal{R}_{S}$ (Moakher, 2002).
This problem may be highly degenerate, even for shapes with a single non-trivial symmetry. For example, consider the bench shape $S$ depicted in Figures LABEL:fig:teaser, 2, LABEL:fig:prop3.1. As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis. One may represent these rotations by the matrices $I$ and $Q:=(-e_{x},e_{y},-e_{z})$, resp., where $e_{x},e_{y},e_{z}$ are the standard basis vectors.
@@
Proposition 3.2 states that one solves the $L^{2}$ regression problem 2 for the bench shape by computing the arithmetic mean of $I$ and $Q$ and then orthogonally projecting this matrix onto $SO(3)$. The arithmetic mean of $I,Q$ is the matrix $M:=(0,e_{y},0)$, and one computes its orthogonal projection onto $SO(3)$ by solving a special Procrustes problem (Gower & Dijksterhuis, 2004). However, the solution to this problem is non-unique, and $f^{*}(S)$ may be any rotation about the $y$-axis, which we illustrate in Figure 4. This shows that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2.
A partial solution.
The previous section shows that solving Equation 1 with the $L^{2}$ loss fails for rotationally-symmetric shapes, which are common in practice. We now present a partial solution to this problem. Suppose we know a finite group ^$\hat{\mathcal{R}}\supseteq\mathcal{R}_{S}$ that contains a shape $S$’s rotational symmetries. We can then quotient the $L^{2}$ loss by $\hat{\mathcal{R}}$ to obtain the following problem:
@@
We prove this proposition in Appendix A.3. In contrast to naïve $L^{2}$ regression, quotient regression learns a function that correctly orients rotationally-symmetric shapes up to a rotation in the group $\hat{\mathcal{R}}$. While this is only a partial solution to the orientation estimation problem, the remainder reduces to a discrete classification problem: Predicting the rotation $Q^{*}\in\hat{\mathcal{R}}$ such that $f^{*}(RS)=RQ^{*}\Omega_{S}$. In the following section, we will show how a solution to this problem allows one to map $RS$ to the canonically-oriented shape $S$.
Recovering an orientation via classification.
@@
By solving the quotient regression problem in Equation 3, one can recover an arbitrarily-rotated shape $RS$’s orientation up to a rotation $Q^{*}\in\hat{\mathcal{R}}$. In this section, we propose training a classifier to predict this rotation $Q^{*}$ given the solution $f^{*}(RS)=RQ^{*}\Omega_{S}$ to the quotient regression problem. We now further assume that the shape $S$’s ground truth orientation $\Omega_{S}$ is the canonical orientation $\Omega_{S}=I$. We show that even if $S$ is symmetric under some group of symmetries $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$, the optimal classifier’s predictions enable one to map $RS$ to the canonically-oriented shape $S$.
Predicting a rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model is an $|\hat{\mathcal{R}}|$-class classification problem. One may train an appropriate classifier by solving the following problem:
$\min_{p_{\phi}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{\begin{subarray}{c}Q\sim U(\hat{\mathcal{R}})\\ S\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\textrm{CE}\left(p_{\phi}(QS),\delta_{Q}\right)\right],$ (5)
@@
where $U(\hat{\mathcal{R}})$ denotes the uniform distribution on $\hat{\mathcal{R}}$, $\textrm{CE}(\cdot)$ denotes the cross-entropy loss, and $\delta_{Q}\in\Delta^{|\hat{\mathcal{R}}|-1}$ is a one-hot vector centered at the index of $Q\in\hat{\mathcal{R}}$. While one may hope that composing the quotient regression model and this classifier yields an orienter that outputs correct orientations $\hat{\Omega}_{RS}=R\Omega_{S}$ regardless of its inputs’ symmetries, recall that Proposition 3.1 shows such an orienter does not exist. However, the following result shows that this pipeline recovers the orientation of a shape $RS$ up to one of its rotational symmetries, which is sufficient for mapping $RS$ to $S$.
Proposition 3.4
@@
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$, and suppose $S$ is canonically-oriented, so $\Omega_{S}=I$. Let $f^{*}:\mathcal{S}\rightarrow SO(3)$ be a solution to Equation 3, so that $f^{*}(RS)=RQ^{*}$ for some $Q^{*}\in\hat{\mathcal{R}}$. Finally, suppose that $p^{*}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}$ solves the following problem:
@@
$\min_{p:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{Q\sim U(\hat{\mathcal{R}})}{\mathbb{E}}\left[\textrm{CE}\left(p(QS),\delta_{Q}\right)\right].$ (6)
@@
Then for any $R\in SO(3)$, $p^{*}(f^{*}(RS)^{\top}RS)$ is the uniform distribution over $\left\{(Q^{*})^{\top}F:F\in\mathcal{R}_{S}\right\}$. For any $(Q^{*})^{\top}F$ in the support of this distribution, $\underbrace{((Q^{*})^{\top}F)^{\top}}_{\textrm{second-stage prediction}}\underbrace{f^{*}(RS)^{\top}}_{\textrm{first-stage prediction}}RS=S$, so using $f^{*}$ and $p^{*}$, one may recover $S$ from the arbitrarily-rotated shape $RS$.
We prove this proposition in Appendix A.4. How does one reconcile this result with Proposition 3.1? The orientation of $(Q^{*})^{\top}F)^{\top}f^{*}(RS)^{\top}RS$ is $F^{\top}\neq I=\Omega_{S}$, so this result does not contradict Proposition 3.1. Rather, it shows that when a shape $S$ is rotationally symmetric, one need only predict the orientation of $RS$ up to one of its symmetries to recover the canonically-oriented $S$. We combine these results in the following section to implement a state-of-the-art method for orientation estimation.

### v2

@@ Abstract
3d orientation, shape analysis, 3d deep learning, geometric deep learning
@@ 1 Introduction
Orientation estimation is a fundamental task in 3D shape analysis which consists of estimating a shape’s orientation axes: its side-, up-, and front-axes. Using this data, one can rotate a shape into canonical orientation, in which the shape’s orientation axes are aligned with the coordinate axes. This task is especially important as a pre-processing step in 3D deep learning, where deep networks are typically trained on datasets of canonically oriented shapes but applied to arbitrarily-oriented shapes at inference time. While data augmentation or equivariant and invariant architectures may improve a model’s robustness to input rotations, these techniques come at the cost of data efficiency and model expressivity (Kuchnik & Smith, 2019; Kim et al., 2023). In contrast, orientation estimation allows one to pre-process shapes at inference time so that their orientation matches a model’s training data.
A shape’s orientation axes are not intrinsic geometric quantities: They are determined by humans based on physical and functional considerations. It is therefore difficult to construct simple geometric algorithms for orientation estimation. However, given a large dataset of canonically oriented shapes, one may pose orientation estimation as a supervised learning problem. This task is challenging, and developing an orientation pipeline that reliably estimates complete orientations of general shapes remains an open problem.
The naïve deep learning approach is to train a model with an $L_{2}$ loss to directly predict a shape’s orientation from a point cloud of surface samples. However, this strategy fails for shapes with rotational symmetries, where the optimal solution to the $L_{2}$ regression problem is the Euclidean mean (Moakher, 2002) of a shape’s orientations over all of its symmetries. In contrast, works such as Poursaeed et al. (2020) discretize the unit sphere into a set of fixed rotations and train a classifier to predict a probability distribution over these rotations, but find that this approach fails for any sufficiently dense discretization of the unit sphere.
@@
Using this strategy, our method accurately estimates a shape’s orientation up to one of its symmetries, which suffices for returning the shape to an upright and front-facing pose. As we are unaware of prior work on this problem, we also benchmark our pipeline’s performance on the well-studied task of up-axis prediction, in which one seeks to return a shape to an upright (but not necessarily front-facing) pose, and find that our work achieves state-of-the-art performance on this task. Unlike previous work, we train and evaluate our model on the entire Shapenet dataset rather than a subset of classes. We further demonstrate its generalization capabilities on Objaverse, a large dataset of real-world 3D models of varying quality.
@@
Our contributions include the following: (1) we identify fundamental obstacles to orientation estimation and study the conditions under which a naïve regression-based approach to orientation estimation fails; (2) we propose a two-stage orientation estimation pipeline that sidesteps these obstacles; (3) we train and test our model on Shapenet and show that it achieves SOTA performance for up-axis prediction and orientation estimation; (4) we use conformal prediction to enable end users to resolve ambiguities in a shape’s orientation; (5) we release our code and model weights to share our work with the ML community.
@@ 2 Related work
Chen et al. (2021) use reinforcement learning to train a model to gradually rotate a shape into upright orientation. While this algorithm performs well, it is evaluated on few classes and is costly to train. Kim et al. (2020) adopt a similar perspective to Fu et al. (2008), but use ConvNets to extract features for a random forest classifier that predicts a natural base. Poursaeed et al. (2020) use orientation estimation as a pretext task to learn features for shape classification and keypoint prediction. They also investigate a pure classification-based approach to orientation estimation that discretizes the 3D rotation group into $K$ rotations and predicts a distribution over these rotations for an arbitrarily-rotated input shape. They find that its performance decays rapidly as $K$ increases, reaching an accuracy as low as 1.6% for $K=100$ rotations.
We also highlight a related literature on canonical alignment. This literature includes works such as Kim et al. (2023); Sajnani et al. (2022); Spezialetti et al. (2020); Zhou et al. (2022), which seek to map arbitrarily-rotated shapes to a class-consistent pose, as well as Katzir et al. (2022); Sun et al. (2021), which seek to learn pose-invariant representations of 3D shapes. These works only attempt to learn a consistent orientation within each class, but this orientation is not consistent across classes and is not generally aligned with the coordinate axes. In contrast, we tackle the more challenging task of inferring a canonical orientation that is consistent across all objects.
More broadly, geometric deep learning studies general methods for exploiting symmetries in data by designing machine learning models that are invariant or equivariant to certain transformations. For instance, Kaba et al. (2023) proposes using learned canonicalization functions to obtain equivariant machine learning methods, and Puny et al. (2022) proposes frame averaging to induce invariance or equivariance in deep learning architectures. Our work may be viewed as an efficient canonicalization method for the specific case of 3D shapes with rotational symmetries.
@@ 3 Method
In this section, we motivate and describe our orientation pipeline. We first identify fundamental obstacles to orientation estimation and show that learning a shape’s orientation with the $L_{2}$ loss fails when the shape is rotationally symmetric. Motivated by these observations, we introduce our two-stage orientation pipeline consisting of a quotient orienter followed by a flipper. Our quotient orienter model solves a regression problem to recover a shape’s orientation up to octahedral symmetries, which commonly occur in real-world shapes. The flipper then predicts one of 24 octahedral flips that returns the first-stage output to canonical orientation. We finally use conformal prediction to enable our flipper to output prediction sets whose size varies with the model’s uncertainty. This allows end users to resolve ambiguities in a shape’s orientation by choosing from a small set of plausible candidate orientations.
@@ 3.1 Orientation estimation under rotational symmetries
In this section, we introduce the orientation estimation problem and motivate our approach. Throughout these preliminaries, we consider 3D shapes $S\in\mathcal{S}$ lying in some space of arbitrary shape representations $\mathcal{S}$. Orientation estimation consists of learning an orienter function $f:\mathcal{S}\rightarrow SO(3)$ that maps a shape $S\in\mathcal{S}$ to a predicted orientation $\hat{\Omega}_{S}\in SO(3)$, where $SO(3)$ denotes the 3D rotation group. An orientation is a rotation matrix $\Omega_{S}$ associated with a shape $S$ that is rotation-equivariant: If one rotates $S$ by $R\in SO(3)$ to obtain $RS$, then $\Omega_{RS}=R\Omega_{S}$. We interpret the columns of $\Omega_{S}=(\omega_{S}^{x},\omega_{S}^{y},\omega_{S}^{z})$ as the side-, up-, and front-axes of $S$, respectively, and say that $S$ is in canonical orientation if $\Omega_{S}=I$. If $S$ is in canonical orientation, then its side-, up-, and front-axes (which we will jointly refer to as its orientation axes) are aligned with the $\{x,y,z\}$ coordinate axes, respectively. We depict a canonically oriented shape $S$ along with its orientation $\Omega_{S}$ in Figure 2.
@@
A shape’s orientation axes are not intrinsic geometric quantities. Instead, they are determined by humans based on physical and functional considerations. For instance, a shape’s up-axis may be determined by how a shape naturally lies under gravity, whereas its front-axis may be determined by the direction from which a human typically interacts with it. It is therefore challenging to devise simple geometric algorithms for orientation estimation.
However, there exist large datasets of shapes such as Shapenet (Chang et al., 2015) that have been assigned a canonical orientation by human annotators. Given a training set $\mathcal{D}$ of shapes $S\in\mathcal{S}$ paired with their ground truth orientations $\Omega_{S}$ (where $\Omega_{S}\equiv I$ for datasets of canonically oriented shapes), a natural strategy for orientation estimation is to parametrize the orienter as a neural network $f_{\theta}$ and solve the following supervised learning problem:
$\min_{f_{\theta}}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\|f_{\theta}(RS)-R\Omega_{S}\|_{F}^{2}\right],$ (1)
where $U(SO(3))$ is the uniform distribution over $SO(3)$.
Many real-world shapes $S$ possess at least one non-trivial rotational symmetry $R$. This rotation leaves the shape unchanged, so $RS=S$. However, since a shape’s orientation is rotation-equivariant, $\Omega_{RS}=R\Omega_{S}\neq\Omega_{S}$, and the same shape $S$ will necessarily be associated with two distinct orientations $\Omega_{S}$ and $R\Omega_{S}$. It follows that the ground truth orienter map $S\mapsto\Omega_{S}$ of a rotationally-symmetric shape is one-to-many and therefore not a function. We formally state and prove this result in Proposition A.1, and depict an instance of this phenomenon in Figure LABEL:fig:prop3.1.
On the other hand, $f_{\theta}$ is a function, which implies that the solution to Equation 1 cannot be the true orienter map. The following proposition shows that for single shapes with rotational symmetries, the solution to Equation 1 is instead Euclidean mean of the rotated orientations $RQ\Omega_{S}$ across all rotations $Q$ in the symmetry group $\mathcal{R}_{S}$ (Moakher, 2002).
Proposition 3.1.
@@
We prove this proposition in Appendix A.2. This problem may be highly degenerate, even for shapes with a single non-trivial symmetry. For example, consider the bench shape $S$ depicted in Figures LABEL:fig:teaser, 2, LABEL:fig:prop3.1. As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis. It is straightforward to show that the solution $f^{*}$ to Equation 2 is non-unique and may be any rotation about the $y$-axis, which we illustrate in Figure 4. (See Appendix A.3 for further details.) This demonstrates that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2, which substantiates the observations of Liu et al. (2016); Poursaeed et al. (2020) that orientation estimation via $L_{2}$ regression typically performs poorly in practice.
@@
3.2 A partial solution.
The previous section shows that orientation estimation via $L_{2}$ regression fails for rotationally-symmetric shapes, which are common in practice. We now present a partial solution to this problem. Suppose we know a finite group ^$\hat{\mathcal{R}}\supseteq\mathcal{R}_{S}$ that contains a shape $S$’s rotational symmetries. We can then quotient the $L_{2}$ loss by $\hat{\mathcal{R}}$ to obtain the following problem:
@@
We prove this proposition in Appendix A.4. In contrast to naïve $L_{2}$ regression, quotient regression learns a function that correctly orients rotationally-symmetric shapes up to a rotation in the group $\hat{\mathcal{R}}$. While this is only a partial solution to the orientation estimation problem, the remainder reduces to a discrete classification problem: Predicting the rotation $Q^{*}\in\hat{\mathcal{R}}$ such that $f^{*}(RS)=RQ^{*}\Omega_{S}$. In the following section, we will show how a solution to this problem allows one to map $RS$ to the canonically oriented shape $S$.
3.3 Recovering an orientation via classification.
@@
By solving the quotient regression problem in Equation 3, one can recover an arbitrarily-rotated shape $RS$’s orientation up to a rotation $Q^{*}\in\hat{\mathcal{R}}$. In this section, we propose training a classifier to predict this rotation $Q^{*}$ given the solution $f^{*}(RS)=RQ^{*}\Omega_{S}$ to the quotient regression problem. We now further assume that the shape $S$’s ground truth orientation $\Omega_{S}$ is the canonical orientation $\Omega_{S}=I$. We show that even if $S$ is symmetric under some group of symmetries $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$, the optimal classifier’s predictions enable one to map $RS$ to the canonically oriented shape $S$.
@@
Predicting a rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model is an $|\hat{\mathcal{R}}|$-class classification problem. While one may hope that composing the quotient regression model and the optimal classifier yields a function that outputs correct orientations $\hat{\Omega}_{RS}=R\Omega_{S}$ regardless of its inputs’ symmetries, such a function cannot exist for symmetric shapes. However, the following result shows that this pipeline recovers the orientation of a shape $RS$ up to one of its rotational symmetries, which suffices for mapping $RS$ to $S$.
Proposition 3.3.
@@
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$, and suppose $S$ is canonically oriented, so $\Omega_{S}=I$. Let $f^{*}:\mathcal{S}\rightarrow SO(3)$ be a solution to Equation 3, so that $f^{*}(RS)=RQ^{*}$ for some $Q^{*}\in\hat{\mathcal{R}}$. Finally, suppose that $p^{*}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}$ solves the following problem:
@@
$\min_{p:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{Q\sim U(\hat{\mathcal{R}})}{\mathbb{E}}\left[\textrm{CE}\left(p(QS),\delta_{Q}\right)\right],$ (5)
@@
where $U(\hat{\mathcal{R}})$ denotes the uniform distribution on $\hat{\mathcal{R}}$, $\textrm{CE}(\cdot)$ denotes the cross-entropy loss, and $\delta_{Q}\in\Delta^{|\hat{\mathcal{R}}|-1}$ is a one-hot vector centered at the index of $Q\in\hat{\mathcal{R}}$. Then for any $R\in SO(3)$, $p^{*}(f^{*}(RS)^{\top}RS)$ is the uniform distribution over $\left\{(Q^{*})^{\top}F:F\in\mathcal{R}_{S}\right\}$. For any $(Q^{*})^{\top}F$ in the support of this distribution, $\underbrace{((Q^{*})^{\top}F)^{\top}}_{\textrm{second-stage prediction}}\underbrace{f^{*}(RS)^{\top}}_{\textrm{first-stage prediction}}RS=S$, so using $f^{*}$ and $p^{*}$, one may recover $S$ from the arbitrarily-rotated shape $RS$.
We prove this proposition in Appendix A.5. The results in this section show that unlike naïve $L_{2}$ regression, a two-stage pipeline consisting of quotient regression followed by discrete classification can successfully recover a shape’s orientation up to its symmetries, provided one quotients the $L_{2}$ objective by a sufficiently large subgroup of $SO(3)$. We combine these results in the following section to implement a state-of-the-art method for orientation estimation.